Circular No. 8786

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 Y1 (KORLEVIĆ-JURIĆ) E. J. Christensen reports his recovery of comet P/1999 DN_{3} (cf. IAUC

7167) on CCD images obtained with the 1.5-m reflector at Mt. Lemmon, with his astrometry provided below. On Dec. 16.4 UT, 60-s images obtained in average seeing showed a moderately condensed 5′′ coma with a 15′′ tail in p.a. 270^{o}. Four co-added 60-s images obtained the following night in poor seeing confirm a faint 10′′ tail in p.a. 260^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Dec. 16.36258  7^{h}31^{m}34.^{s}32  +45^{o}48′08′′.8  20.0
16.36796  7 31 34.06  +45 48 10.0
16.37348  7 31 33.81  +45 48 10.9
16.37911  7 31 33.60  +45 48 12.2
17.28139  7 30 55.98  +45 51 21.3  20.7
17.28505  7 30 55.83  +45 51 21.4  20.5
17.28874  7 30 55.69  +45 51 22.2  20.5
17.29240  7 30 55.58  +45 51 23.1  20.7

The indicated correction to the prediction on MPC 54168 is ΔT = −2.0 days. The following linked orbital elements are by B. G. Marsden:

Epoch = 1998 Sept. 24.0 TT

T = 1998 Sept.30.4678 TT  ω = 161.^{o}2491
e = 0.135043  Ω = 5.9549 }2000.0
q = 3.907590 AU  i = 18.7192

a = 4.517671 AU n^{o} = 0.1026437 P = 9.60 years Epoch = 2008 May 14.0 TT

T = 2008 May 8.7944 TT  ω = 161.^{o}6439
e = 0.135853  Ω = 5.8430 }2000.0
q = 3.893755 AU  i = 18.7301

a = 4.505892 AU n^{o} = 0.1030464 P = 9.56 years

COMET C/2003 B3 (SOHO) Another Kreutz sungrazing comet found on SOHO website images (cf.

IAUC 8783), being of stellar appearance (mag ∼ 7.5–8):

Comet 2003 UT α_{2000} δ_{2000} Inst. F MPEC
C/2003 B3 Jan. 22.696 20^{h}32.^{m}4 −21^{o}22′ C3 HS 2006-X25
2006 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green